OpenStudy (anonymous):

A box contains 6 white balls, 4 black balls, and 1 red ball. How many red balls must be added to the box so that the probability of drawing a red ball is 8/9? HOw many black balls must be added to the original box so that the probability of drawing a white ball is 1/9?

OpenStudy (cwrw238):

the total number of balls after adding more red balls must be a multiple of 9 ( for the probability to be 8/9) then the probability of drawing a non red ball will be 1/9 - and there will be 10 of these. so can you figure out how many reds to add now?

OpenStudy (cwrw238):

what will be the total number of balls after the red balls have been added?

OpenStudy (anonymous):

I am so very confused with this question. I have always had a problem with these :) I am trying to figure this out with what u said.

OpenStudy (cwrw238):

hint - there are 10 white and black balls - and that is 1/9 of the total (since probability of picking one is 1/9)

OpenStudy (zarkon):

solve \[\frac{1+x}{6+4+1+x}=\frac{8}{9}\] for \(x\)